# Thermalizer: Stable autoregressive neural emulation of spatiotemporal chaos

### v1

@@ Abstract
Autoregressive surrogate models (or emulators) of spatiotemporal systems provide an avenue for fast, approximate predictions, with broad applications across science and engineering. At inference time, however, these models are generally unable to provide predictions over long time rollouts due to accumulation of errors leading to diverging trajectories. In essence, emulators operate out of distribution, and controlling the online distribution quickly becomes intractable in large-scale settings. To address this fundamental issue, and focusing on time-stationary systems admitting an invariant measure, we leverage diffusion models to obtain an implicit estimator of the score of this invariant measure. We show that this model of the score function can be used to stabilize autoregressive emulator rollouts by applying on-the-fly denoising during inference, a process we call thermalization. Thermalizing an emulator rollout is shown to extend the time horizon of stable predictions by an order of magnitude in complex systems exhibiting turbulent and chaotic behavior, opening up a novel application of diffusion models in the context of neural emulation.
@@ 3 Related work
Diffusion models for PDEs: In (Lippe et al., 2023), an autoregressive model is used to predict the next timestep, and then a small number of “refinement” steps are applied to the prediction. These refinement steps are trained on a denoising loss, and can therefore be interpreted as DDPM denoising steps, analogous to our thermalization steps. Therefore our work is similar in nature, however with the key differences that their denoiser is a conditional model, and that our denoising model is constructed separately to the emulator and applied adaptively at inference time. Additionally, we investigate rollouts over many more autoregressive passes, up to $1e^{5}$ steps.
Diffusion models have also been applied to generating realizations of turbulent fields (Lienen et al., 2023), and emulation in the form of autoregressive conditional models (Kohl et al., 2023; Price et al., 2023), where the probabilistic nature of predictions is able to improve modeling of chaotic dynamics. To reduce the computational cost of generating next step predictions, (Gao et al., 2024) performs the conditional generation in a learned, low-dimensional latent space. Diffusion models have also been used for statistical downscaling in atmospheric weather modelling (Mardani et al., 2023; Wan et al., 2023).
@@ 4.1 Emulator
Figure 4: Mean squared error between neural network trajectories and numerical model trajectories for different baseline architectures (Unet and Dilated ResNet), starting from the same initial conditions. We show $40$ Kolmogorov flow trajectories in the top two panels, and $20$ QG flows in the lower two panels. Thermalized flows remain around a constant MSE long after non-thermalized flows have become unstable. On the right, we show histograms for the total number of thermalization steps along the full trajectories.
@@ 4.2 Thermalizer
Thus, assuming that the noise level is correctly estimated, and an accurate denoising model, the thermalizer step contracts towards equilibrium. Moreover, for nearly Gaussian perturbations, the LHS can be substantially smaller than the original error. It is important to emphasize that even approximate transports can go a long way to stabilize rollouts, as long as they can mitigate the distribution shift. Figure 5 illustrates the principle of the thermalizer.
@@ 5.2 Quasi-geostrophic turbulence
In Figure 4, we show the MSE with respect to the numerical model, for different neural network models and flow configurations. We observe that initially the MSE increases as trajectories diverge over the first few hundred steps, and settle to approximately unity MSE. In this regime, the state vectors are stable, but decorrelated. States with MSE$>>1$ are unstable. Over $\approx 2000$ timesteps emulator trajectories all go unstable (with the exception of the Kolmogorov DRN, where many trajectories stay stable for much longer). However all thermalized trajectories stay stable for the full $10,000$ steps, due to the corrective steps applied during the rollout.
In the right panels, we show the number of thermalization steps applied along the entire trajectory for all models and flow configurations. In the case of the QG Unet, the thermalization is highly efficient, with just $\sim 140$ thermalization steps required on average to stabilise each $10,000$ step trajectory. In terms of computational cost, we run a forward pass of the noise-classifying head of the thermalizer at every timestep, which is $60\%$ the cost of a Unet emulator step, as we are only activating the downsampling and noise-classifying components of the equivalent-sized thermalizer Unet. Additional forward passes of the full thermalizer network are only run in the case where $s_{\mathrm{pred}}>s_{\mathrm{init}}$, so given an accurate noise classifier, the algorithm is inherently efficient and only applies corrective denoising steps when necessary.
@@ Acknowledgements
The authors would like to thank Fabrizio Falasca, Pavel Perezhogin, Stephane Mallat, Etienne Lempereur, Rudy Morel, Freddy Bouchet, Eric Vanden-Eijnden and Edouard Oyallon for many valuable discussions. This project is supported by Schmidt Sciences, LLC.
@@ Appendix C Diffusion model thermalizer
To implement the thermalizer as a diffusion model, we use the same core architecture as the Unet used for the emulator. As discussed in section 4.2, a key component of the thermalizer is some prediction of the noise level. To augment this standard Unet implementation with a noise-classifying head, we take the lowest-dimensional representation from the Unet, and add an additional $2$ convolutional layers, with kernel size of $3$ and stride of $1$. We then vectorize the output of the last convolutional filter into a length $4096$ vector. This is then passed through 2 more linear layers of size $1000$. This final vector of length $1000$ represents the predicted categorical distribution over noise levels for the input image. The total number of learnable parameters for this network is $53$M. We implement a method to run a forward pass through the noise-classifying component of the network only, for computational efficiency during rollouts. Also note that since our network does not include the noise level as an input scalar, our diffusion model does not need have additional embedding components.
To optimise the network, equation 8 is broken up into mini-batches of size $64$. Again we use the AdamW optimizer, with a learning rate of $2e^{-5}$, and train the thermalizer for $35$ epochs. Given that at inference time, we generally only use low noise levels of approximately $s\lesssim 20$, we experimented with training the model only on these low noise levels. However we found that performance degraded significantly, so sample noise levels $s$ uniformly across the full range during training. Noise levels are set using a cosine variance scheduler. All model architecture, training and inference codes will be made publicly available via github on publication.
In terms of the best performing thermalization criteria, we found $s_{\mathrm{init}}=7$ and $s_{\mathrm{stop}}=4$ optimal for the Kolmogorov Unet, $s_{\mathrm{init}}=10$ and $s_{\mathrm{stop}}=5$ optimal for the Kolmogorov DRN. For QG we use $s_{\mathrm{init}}=9$ and $s_{\mathrm{stop}}=4$ for the Unet, and $s_{\mathrm{init}}=6$ and $s_{\mathrm{stop}}=3$ for the DRN emulator.
@@ Appendix D Inspecting a single thermalization step
We can then continue running the emulator on the snapshots with (middle row) and without (top row) applying this single thermalization step, and observe that the eventual instability develops sooner in the non-thermalized (top) row. This can be seen both visually in the potential vorticity fields as accumulation of high-frequency signal, and the growing field magnitudes shown by the colorbars. This experiment illustrates the effect that the subtle thermalization steps have on the stability of rollouts. The same procedure is repeated for other randomly selected snapshots in Figures 10 and 11.
@@
Figure 11: Same as Figures 9 and 10, except with a different initial condition.
Appendix E Additional results
Here we show the vorticity fields for $25$ random samples from the $40$ trajectories shown in Figure 1, at the end of the $10,000$ steps. In In Figure 15, we show results for the numerical model. Above each panel, we show the predicted noise level for each flow state. In Figures 16 and 17, we show the equivalent for the emulator and thermalized trajectories. Note that in the emulator case, after $10,000$ steps the fields have gone so far out of distribution that the predicted noise levels are unreliable.
In Figures 18, 19 and 20, we show the upper layer potential vorticity for all $20$ QG trajectories after $10,000$ steps, for the numerical, emulator, and thermalized models respectively.

### v2

@@ Abstract
Autoregressive surrogate models (or emulators) of spatiotemporal systems provide an avenue for fast, approximate predictions, with broad applications across science and engineering. At inference time, however, these models are generally unable to provide predictions over long time rollouts due to accumulation of errors leading to diverging trajectories. In essence, emulators operate out of distribution, and controlling the online distribution quickly becomes intractable in large-scale settings. To address this fundamental issue, and focusing on time-stationary systems admitting an invariant measure, we leverage diffusion models to obtain an implicit estimator of the score of this invariant measure. We show that this model of the score function can be used to stabilize autoregressive emulator rollouts by applying on-the-fly denoising during inference, a process we call thermalization. Thermalizing an emulator rollout is shown to extend the time horizon of stable predictions by two orders of magnitude in complex systems exhibiting turbulent and chaotic behavior, opening up a novel application of diffusion models in the context of neural emulation.
@@ 3 Related work
Diffusion models for PDEs: In (Lippe et al., 2023), an autoregressive model is used to predict the next timestep, and then a small number of “refinement” steps are applied to the prediction. These refinement steps are trained on a denoising loss, and can therefore be interpreted as DDPM denoising steps, analogous to our thermalization steps. Therefore our work is similar in nature, however with the key differences that their denoiser is a conditional model, and that our denoising model is constructed separately to the emulator and applied adaptively at inference time. Additionally, we investigate rollouts over many more autoregressive passes, up to $1e^{5}$ steps. Diffusion models have also been applied to generating realizations of turbulent fields (Lienen et al., 2023), and emulation in the form of autoregressive conditional models (Kohl et al., 2023; Price et al., 2023), where the probabilistic nature of predictions is able to improve modeling of chaotic dynamics. To reduce the computational cost of generating next step predictions, (Gao et al., 2024) performs the conditional generation in a learned, low-dimensional latent space, and (Shehata et al., 2025) develops a more efficient conditional generation algorithm. Diffusion models have also been applied to large-scale weather modelling, both for the purposes of emulation (Price et al., 2023; Rühling Cachay et al., 2024) and for statistical downscaling (Mardani et al., 2023; Wan et al., 2023).
@@ 4.1 Emulator
Figure 4: Mean squared error between neural network trajectories and numerical model trajectories for different baseline architectures (Unet and Dilated ResNet), starting from the same initial conditions. We show $40$ Kolmogorov flow trajectories in the top two panels, and $20$ QG flows in the lower two panels. Thermalized flows remain around a constant MSE long after non-thermalized flows have become unstable. The region where state vectors are decorrelated in time, but stable, is shown in shaded pink. On the right, we show histograms for the total number of thermalization steps along the full trajectories.
@@ 4.2 Thermalizer
Thus, assuming that the noise level is correctly estimated, and an accurate denoising model, the thermalizer step contracts towards equilibrium. Moreover, for nearly Gaussian perturbations, the LHS can be substantially smaller than the original error. It is important to emphasize that even approximate transports can go a long way to stabilize rollouts, as long as they can mitigate the distribution shift.
@@ 5.2 Quasi-geostrophic turbulence
In Figure 4, we show the MSE with respect to the numerical model, for different neural network models and flow configurations. We observe that initially the MSE increases as trajectories diverge over the first few hundred steps, and settle to approximately unity MSE. In this regime, shaded in pink, the state vectors are stable but decorrelated. States with MSE$>>1$ are unstable. Over $\approx 2000$ timesteps emulator trajectories all go unstable (with the exception of the Kolmogorov DRN, where many trajectories stay stable for much longer). However all thermalized trajectories stay stable for the full $10,000$ steps, due to the corrective steps applied during the rollout. In the right panels, we show the number of thermalization steps applied along the entire trajectory for all models and flow configurations. In the case of the QG Unet, the thermalization is highly efficient, with just $\sim 140$ thermalization steps required on average to stabilise each $10,000$ step trajectory. In terms of computational cost, we run a forward pass of the noise-classifying head of the thermalizer at every timestep, which is $60\%$ the cost of a Unet emulator step, as we are only activating the downsampling and noise-classifying components of the equivalent-sized thermalizer Unet. Additional forward passes of the full thermalizer network are only run in the case where $s_{\mathrm{pred}}>s_{\mathrm{init}}$, so given an accurate noise classifier, the algorithm is inherently efficient and only applies corrective denoising steps when necessary.
@@ Acknowledgements
The authors would like to thank Fabrizio Falasca, Pavel Perezhogin, Stephane Mallat, Etienne Lempereur, Freddy Bouchet, Eric Vanden-Eijnden and Edouard Oyallon for many valuable discussions. This project is supported by Schmidt Sciences, LLC.
@@ Appendix C Diffusion model thermalizer
C.1 Architecture
To implement the thermalizer as a diffusion model, we use the same core architecture as the Unet used for the emulator. As discussed in section 4.2, a key component of the thermalizer is some prediction of the noise level. To augment this standard Unet implementation with a noise-classifying head, we take the lowest-dimensional representation from the Unet, and add an additional $2$ convolutional layers, with kernel size of $3$ and stride of $1$. We then vectorize the output of the last convolutional filter into a length $4096$ vector. This is then passed through 2 more linear layers of size $1000$. This final vector of length $1000$ represents the predicted categorical distribution over noise levels for the input image. The total number of learnable parameters for this network is $53$M. We implement a method to run a forward pass through the noise-classifying component of the network only, for computational efficiency during rollouts. Also note that since our network does not include the noise level as an input scalar, our diffusion model does not need have additional timestep embedding components.
C.2 Optimization
To optimise the network, equation 8 is broken up into mini-batches of size $64$. Again we use the AdamW optimizer, with a learning rate of $2e^{-5}$, and train the thermalizer for $35$ epochs. Given that at inference time, we generally only use low noise levels of approximately $s\lesssim 20$, we experimented with training the model only on these low noise levels. However we found that performance degraded significantly, so sample noise levels $s$ uniformly across the full range during training. Noise levels are set using a cosine variance scheduler. All model architecture, training and inference codes can be found at https://github.com/Chris-Pedersen/thermalizer.
C.3 Thermalizer settings: $s_{\mathrm{init}}$ and $s_{\mathrm{stop}}$
A key component of the algorithm is the adaptive nature of thermalization - which allows for stabilisation with a minimal interruption of temporal dynamics. There are two free parameters here - $s_{\mathrm{init}}$ which determines the estimated noise level at which the thermalizer starts to apply corrections to the state vector, and $s_{\mathrm{stop}}$, which sets the lowest noise level to which we run the denoising process. To find the optimum settings, we run a simple grid search on $s_{\mathrm{init}}\in[10,6]$ and $s_{\mathrm{stop}}\in[5,2]$ (note that we need $s_{\mathrm{start}}>s_{\mathrm{stop}}$. For each pair of $s_{\mathrm{init}}$ and $s_{\mathrm{stop}}$, we run a thermalized trajectory for $41e$ 4 steps. We chose the combination of $s_{\mathrm{init}}$ and $s_{\mathrm{stop}}$ in which the kinetic energy spectra of the thermalized trajectories best matches the kinetic energy in the numerical model. This experiment is repeated independently for each flow configuration, and each emulator - so this procedure was run a total of $4$ times. We found $s_{\mathrm{init}}=7$ and $s_{\mathrm{stop}}=4$ optimal for the Kolmogorov Unet, $s_{\mathrm{init}}=10$ and $s_{\mathrm{stop}}=5$ optimal for the Kolmogorov DRN. For QG we use $s_{\mathrm{init}}=9$ and $s_{\mathrm{stop}}=4$ for the Unet, and $s_{\mathrm{init}}=6$ and $s_{\mathrm{stop}}=3$ for the DRN emulator.
@@ Appendix D Inspecting a single thermalization step
We can then continue running the emulator on the snapshots with (middle row) and without (top row) applying this single thermalization step, and observe that the eventual instability develops sooner in the non-thermalized (top) row. This can be seen both visually in the potential vorticity fields as accumulation of high-frequency signal, and the growing field magnitudes shown by the colorbars. This experiment illustrates the effect that the subtle thermalization steps have on the stability of rollouts. The same procedure is repeated for other randomly selected snapshots in Figures 9 and 10.
@@
Figure 10: Same as Figures 8 and 9, except with a different initial condition.
Appendix E Additional results and figures
In Figure 11 we show a schematic illustrating the principle of the thermalizer. In Figure 12, we show $s_{\mathrm{pred}}$ along the full $10000$ step trajectory, for the Unet emulator in both Kolmogorov and QG flows. In gray lines we show $s_{\mathrm{pred}}$ for the emulator trajectories, thermalized trajectories are shown in blue, and the numerical model in red. Initially, $s_{\mathrm{pred}}$ is low for all three different models, as the state vector has not had a chance to accumulate significant error. We see that $s_{\mathrm{pred}}$ increases as the un-thermalized trajectories accumulate error - eventually going unstable, at which point our noise-classifying output becomes unreliable, leading to the vertical gray lines. In contrast, the blue lines are constrained to stay at a low noise level, never exceeding $s_{\mathrm{pred}}=16$. This is the consequence of the thermalizer keeping states in-distribution. Finally the red lines for the numerical model are shown as a sanity check that our noise-classifying head is able to consistently identify that the numerical model states are all in-distribution.
In Figure 13, we inspect the number of thermalization steps along a $500$ step segment of the trajectories show in Figures 1 and 15. For each of the 40 trajectories, and each of the $500$ steps between timesteps $5,000$ and $5,500$, we show the number of thermalization steps applied to the state vector. We see that the amount of thermalization applies varies strongly along a trajectory, with steps ranging from 00 to $16$. This indicates that the flow oscillates between periods of no correction, and periods where significant correction is applied.
Next we show the vorticity fields for $25$ random samples from the $40$ trajectories shown in Figure 1, at the end of the $10,000$ steps. In In Figure 16, we show results for the numerical model. Above each panel, we show the predicted noise level for each flow state. In Figures 17 and 18, we show the equivalent for the emulator and thermalized trajectories. Note that in the emulator case, after $10,000$ steps the fields have gone so far out of distribution that the predicted noise levels are unreliable.
In Figures 19, 20 and 21, we show the upper layer potential vorticity for all $20$ QG trajectories after $10,000$ steps, for the numerical, emulator, and thermalized models respectively.
@@ Appendix E Additional results and figures
Figure 13: Number of thermalization steps along thermalized trajectories. We inspect the number of thermalization steps for each of the $40$ trajectories at each timestep, between timestep $5000$ and $5500$, for both the Unet and DRN emulators on Kolmogorov flow.